OpenStudy (anonymous):

Can anyone help me a car travels 28 miles north and then turns east. How far did the car travel east if the destination is 53 miles from the starting point

OpenStudy (anonymous):

You subtract 28 from 53.

OpenStudy (anonymous):

Ok so that going to be -25 right

OpenStudy (anonymous):

a^2+b^2=c^2 a=28 b=wat your trying to find and c=53 just plug in and solve

OpenStudy (anonymous):

Yeah. ^^ thats what im trying to say.

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

so you travel 45 miles east

OpenStudy (anonymous):

So the answer would be -2027

OpenStudy (anonymous):

No the answer would be 25 miles east.

OpenStudy (anonymous):

28^2+b^2=53^2 784+b^2=2809 b^2=2025 square root it and then you get 45 miles east

OpenStudy (anonymous):

Because you take 53 and subtract 28 and that gives you 25.

OpenStudy (anonymous):

Ok now i got it thx

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

If the destination is 53 miles away and you already traveled 28 you'll have 25 miles to be at yor destination.

OpenStudy (anonymous):

Yeah your so right thx again

OpenStudy (anonymous):

No problem. (: